HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

BOUND-BOUND AND BOUND-FREE SPECTROSCOPY OF COLD TRAPPED NEGATIVE IONS

\underline{R. WESTER} , Institut für Ionenphysik und Angewandte Physik, Universität Innbruck, 6020 Innsbruck, Austria

Cryogenic radiofrequency ion traps are a versatile tool to prepare molecular ions at cold translational as well as rovibrational temperatures ^{1}^{,}^{2}. For negative ions, electron affinities and internal state populations can then be studied using photodetachment spectroscopy near threshold. We have used this approach to investigate rotational cooling of OH^{−} anions ^{3}^{4} and to determine the electron affinities of CN, C_{3}N, and C_{2} with better accuracies^{5}^{,}^{6}^{,}^{7}. In the case of C_{3}N^{−} this provided evidence for a dipole-bound state just below threshold, in good agreement with theoretical calculations. Furthermore, we have studied high-resolution bound-bound spectroscopy of a vibrational overtone in OH^{−} ^{8} and an electronic transition in C_{2}^{−} ^{9}, which provided direct access to the translational velocity distributions of the ions. Recently, we have also observed the fluorescence from the excited B-state in C_{2}^{−}, mediated by vibrational quenching collisions^{10}.

1D. Gerlich, Phys. Scr. T59, 256 (1995)

2R. Wester, J. Phys. B 42, 154001 (2009)

3D. Hauser, S. Lee, F. Carelli, S. Spieler, O. Lakhmanskaya, E. S. Endres, S. S. Kumar, F. Gianturco, R. Wester, Nat. Phys. 11, 467 (2015)

4E. S. Endres, G. Egger, S. Lee, O. Lakhmanskaya, M. Simpson, R. Wester, J. Mol. Spectrosc. 332, 134 (2017)

5M. Simpson, M. Nötzold, A. Schmidt-May, T. Michaelsen, B. Bastian, J. Meyer, R. Wild, F. A.

Gianturco, M. Milovanovic;, V. Kokoouline, R. Wester, J. Chem. Phys. 153, 184309 (2020)

6M. Simpson, M. Nötzold, T. Michaelsen, R. Wild, F. A. Gianturco, R. Wester, Phys. Rev. Lett.

127, 043001 (2021)

7S. Purushu et al., (in preparation)

8O. Lakhmanskaya, M. Simpson, R. Wester, Phys. Rev. A 102, 012809 (2020)

9M. Nötzold, R. Wild, C. Lochmann, R. Wester, Phys. Rev. A 106, 023111 (2022)

10M. Nötzold et al., (in preparation)

p-number: p244  Submitted on Fri, 23 Jun 2023 17:13:53 +0200